A student write three consecutive odd numbers, he multiply first by 2, second number by 3 and third one by 5, the sum of all 6 numbers is 400, which is lowest odd number?


Answer by mananth(16949)

let the odd integers be x,x+2,x+4
2x,3(x+2), 5(x+4)
x+x+2+x+4+2x+3(x+2)+5(x+4)=400
x+x+2+x+4+2x+3x+6+5x+20=400

13x=400-36
13x=364


solve for x